(2) Pions -ons) are subatomic particles with a half-life of 2.6 x 10 seconds as measured by a clock at rest with respect to them. This means that after the passage of this amount of this time as measured by a clock traveling with them, half of them will have decayed. A batch of 50,000 pions with B-0.92 is created during an experiment at CERN. They are directed along a vacuum line to a detector 50 meters away. How many will arrive at the detector? (5 points) HINT: the two events (in the stationary lab frame) are: (i) creation of the pions atx-0,and (ii) their arrival at the detector at x-50 meters away, while they travel at -092c. The number of particles which still survive after N half- lives is 1/2N of the original number.
Answer #1

Wrt stationary lab frame

No. Of half lives= 50/(0.92*3*108*2.6*10-8) = 7 approx

No of surviving particle =50000/27=390.625 pions.